2 printers can print 4 reams of paper in 30 min. If you had 3 printers and 5 reams to print, how
many minutes would it take?

Answer by stanbon(75887)

2 printers can print 4 reams of paper in 30 min. If you had 3 printers and 5 reams to print, how many minutes would it take?
----
time is directly related to # of reams and indirectly related to # of printers
-----
t = k*r/p
-----
Solve for "k" using " 2 printers can print 4 reams of paper in 30 min".
30 = k*4/2
----
k = 15
-----
Equation:
t = 15*r/p
-----
Question:: If you had 3 printers and 5 reams to print,
how many minutes would it take?
time = 15*5/3
time = 25 minutes
-----------
Cheers,
Stan H.
----------------